# ContractSkill: Repairable Contract-Based Skills for Multimodal Web Agents

### 20 Mar 2026 v1

@@ 6.2. MiniWoB
6.3. Cross-Model Transfer
@@
Table 5. Cross-model transfer results compared against the target model’s own self-generated skill baseline. Aggregate rows merge Layer 1 and Layer 2 within each benchmark.
@@ 7.2. Template-Level MiniWoB Patterns
Figure 3 complements the aggregate MiniWoB ablation with template-level pass rates. Many easy templates are already saturated across all methods, so the overall gain does not come from uniform improvement everywhere. Instead, the advantage of full ContractSkill concentrates on a smaller set of templates that require reliable continuation after partial progress. The clearest gains appear on the two login templates and on mw_m3_click_menu_2, where the repaired artifact clearly outperforms both skill_no_repair and text_only_rewrite.
The same figure also clarifies the remaining boundary cases. Some templates, such as mw_m1_click_link and mw_m2_enter_password, remain unsolved for every method. A different failure mode appears on mw_m4_click_collapsible, where no-skill execution outperforms all skill-based variants. This pattern is consistent with the mechanism suggested by Table 6. Structured repair helps most when there is a reusable procedural skeleton to preserve and patch, but it is not a universal substitute for exploration on every template.
@@
As a complementary VWA analysis, Figure 4 summarizes the residual failure structure across both source models. Relative to No-Skill, ContractSkill leaves fewer failures concentrated in broad late-stage profiles, especially step-6+ outcomes and loop-timeout-heavy mixes, and exposes more failures as localized selector or assertion errors. This shift does not remove all failures, but it makes the remaining ones more structured and more informative for targeted repair.
Figure 4. Supplementary VWA failure analysis across Qwen3.5-Plus and GLM-4.6V. The top panel shows verifier error codes among failed episodes. The bottom panel shows the first failing step bucket. Across both models, ContractSkill reduces the dominance of late coarse failures that are common in No-Skill execution and exposes a larger share of residual failures as earlier, more localizable breakdowns.
Two supplementary stacked horizontal bar charts for VisualWebArena arranged vertically. The top chart shows failure-code distributions for no-skill, self-generated skill, and ContractSkill under Qwen3.5-Plus and GLM-4.6V. The bottom chart shows the corresponding first-failing-step distributions. ContractSkill shifts failures away from late coarse outcomes and toward more localized failure categories.
@@ 8. Discussion and Limitations
At the same time, our current scope is intentionally narrow. We study repair for single multimodal web skill artifacts rather than a full lifelong skill library. The artifact schema is tailored to browser interaction and does not yet cover the broader action space of desktop environments such as OSWorld or Agent S (Xie et al., 2024; Agashe et al., 2024). Our transfer experiments test cross-model reuse within the same benchmark family, which is a meaningful but still limited notion of portability. In addition, deterministic verification is strongest when success can be expressed through programmatic checks over URL, DOM state, or form values; visually ambiguous or highly dynamic web pages remain harder to verify reliably. Finally, the released bundles are not uniformly instrumented for deeper analyses such as stability curves, patch-size distributions, or error-code breakdowns. For this reason, we center the paper on evidence that is consistently available across all settings, namely success, interaction cost, and cross-model reuse. We view richer longitudinal skill-bank analysis as a natural next step rather than a prerequisite for the present claim.
@@ 9. Conclusion
This paper examines a central weakness of multimodal web agents. AI-generated skills are easy to produce, but often too fragile to trust, reuse, or transfer. We present ContractSkill, a framework that turns a draft textual skill into a contracted executable artifact and improves it through deterministic verification, step-level fault localization, and minimal patch repair. Across both VisualWebArena and MiniWoB, ContractSkill consistently outperforms self-generated skills, while the repaired artifacts remain reusable across GLM-4.6V and Qwen3.5-Plus. Taken together, these results suggest a broader shift in how skills should be treated in web agents. Skills should not be disposable prompts that vanish after one run. They should instead be external procedural objects that can be audited, repaired, accumulated, and shared across models. We hope this perspective helps move multimodal web agents from one-shot prompting toward maintainable skill libraries.

### 15 Apr 2026 v3

@@ 6.2. MiniWoB
6.3. Portability Under Matched Transfer Layers
@@
Table 5. Portability under matched transfer layers compared against the target model’s own self-generated skill baseline. Aggregate rows merge Layer 1 and Layer 2 within each benchmark.
@@ 7.2. Template-Level MiniWoB Patterns
Figure 3 adds template-level pass rates to the aggregate MiniWoB ablation. Many easy templates are already saturated across all methods, so the overall gain does not come from uniform improvement everywhere. Instead, the advantage of full ContractSkill concentrates on a smaller set of templates that require reliable continuation after partial progress. The clearest gains appear on the two login templates and on mw_m3_click_menu_2, where the repaired artifact clearly outperforms both skill_no_repair and text_only_rewrite.
The same figure also clarifies the remaining boundary cases. Some templates remain unsolved for every method, including mw_m1_click_link and mw_m2_enter_password. A different failure mode appears on mw_m4_click_collapsible, where no-skill execution outperforms all skill-based variants. This pattern is consistent with the mechanism suggested by Table 6. Structured repair helps most when there is a reusable procedural skeleton to preserve and patch, but it is not a universal substitute for exploration on every template.
@@
(a) Verifier error codes among failed episodes.
(b) First failing step bucket among failed episodes.
Figure 4. VWA failure analysis across Qwen3.5-Plus and GLM-4.6V. Across both models, ContractSkill reduces the dominance of late coarse failures that are common in No-Skill execution and exposes a larger share of residual failures as earlier, more localizable breakdowns.
Two stacked horizontal bar charts for VisualWebArena arranged vertically. The top chart shows failure-code distributions for no-skill, self-generated skill, and ContractSkill under Qwen3.5-Plus and GLM-4.6V. The bottom chart shows the corresponding first-failing-step distributions. ContractSkill shifts failures away from late coarse outcomes and toward more localized failure categories.
Figure 4 summarizes the residual failure structure across both source models in VisualWebArena. Panel (a) shows verifier error codes and Panel (b) shows the first failing step bucket. Relative to No-Skill, ContractSkill leaves fewer failures concentrated in broad late-stage profiles, especially step-6+ outcomes and loop-timeout-heavy mixes, and exposes more failures as localized selector or assertion errors. Manual inspection suggests that self-generated skills commonly fail through early wrong selectors, missing preconditions, and premature stop conditions. These committed but brittle procedures often truncate exploration before the agent can recover, which helps explain why Self-Generated Skill can trail direct acting while still using fewer steps and tokens. This shift does not remove all failures, but it makes the remaining ones more structured and more informative for targeted repair.
@@ 8. Discussion and Limitations
At the same time, our current scope is intentionally narrow. We study repair for single multimodal web skill artifacts rather than a full lifelong skill library. The artifact schema is tailored to browser interaction and does not yet cover the broader action space of desktop environments such as OSWorld or Agent S (Xie et al., 2024; Agashe et al., 2025). Our transfer experiments test cross-model reuse within the same benchmark family, which is a meaningful but still limited notion of portability. In addition, deterministic verification is strongest when success can be expressed through programmatic checks over URL, DOM state, or form values visually ambiguous or highly dynamic web pages remain harder to verify reliably. Finally, the released bundles are not uniformly instrumented for deeper analyses such as stability curves, patch-size distributions, or error-code breakdowns. For this reason, we center the paper on evidence that is consistently available across all settings, namely success, interaction cost, and cross-model reuse. We view richer longitudinal skill-bank analysis as a natural next step rather than a prerequisite for the present claim.
@@ 9. Conclusion
In this paper, we present ContractSkill, a framework that turns a draft textual skill into a contracted executable artifact and improves it through deterministic verification, step-level fault localization, and minimal patch repair. Across both VisualWebArena and MiniWoB, ContractSkill consistently outperforms self-generated skills, while the repaired artifacts remain reusable across GLM-4.6V and Qwen3.5-Plus. Taken together, these results suggest that skills in web agents should be treated as external procedural objects that can be audited, repaired, accumulated, and shared across models. We hope this perspective helps move multimodal web agents from one-shot prompting toward maintainable skill libraries.